Solve by substitution.
3x-5y=4
y=4-2x

Answer by jim_thompson5910(35256)

3x-5y=4

3x-5(4-2x)=4 ... plug in y = 4 - 2x

3x-20+10x=4

13x-20=4

I'll let you finish. You need to solve for x. Once you find x, use it and plug it into y = 4 - 2x to find y.